Suppose the demand for soccer tickets is given by: Q=1000−20PQ=1000−20P. What is the consumer surplus when price is $20 per ticket? Write the dollar amount?
Answer #1

Answer

Q=1000-20P

P=20

Q=1000-20*20=600

the quantity demanded is 600 units

converting the demand curve to the inverse demand curve to find y-axis intercept of the curve which is also called consumer reservation price

Q=1000-20P

20P=1000-Q

P=50-(1/20)Q

consumer reservation price =50

CS=0.5*(consumer reservation price-P)*Q

=0.5*(50-20)*600

=$9000

the consumer surplus is $9000